OpenStudy (anonymous):

Experiment with Gallium: The goal of my experiment is to prove that the heat capacity of gallium is 0.37 J/gC by choosing masses and initial temperatures of both water and gallium. I'm solving for the final temperature using this equation: Mw Cw ΔT = Mg Cg ΔT, where M is mass, C is heat capacity, and ΔT is (final temperature - initial temperature). w refers to water and g refers to gallium. I chose a cold temperature for water and a warmer temperature for gallium, both below the melting point of gallium, but each time the temperature of the mixture is cooler than the water temperature.

OpenStudy (anonymous):

Usually when a cold temperature is mixed with a warmer temperature, the resulting mixture has a temperature than is in the middle. Does this concept not apply to gallium?

OpenStudy (anonymous):

It does, so it's odd that you'd measure that...

OpenStudy (xishem):

Go ahead and show your calculations. Your procedure seems okay, so the error may lie in how you're substituting in your values.

OpenStudy (unklerhaukus):

what temperature is your gallium gonna be?

OpenStudy (anonymous):

Mw = 14.333g Cw = 4.18 J/gC Initial Temp. of Water = 22 C Mg = 21.217 g Cg = 0.37 J/gC Initial Temp. of Gallium = 25 C Mw Cw ΔT = Mg Cg ΔT (14.333g)(4.18 J/gC)(Final Temp. - 22C) = (21.217g)(0.37 J/gC)(Final Temp. - 25C) (59.912 J/C)(Final Temp. - 22C) = (7.850 J/C)(Final Temp.-25 C) -- I distributed here -- (59.912 J/C * Final Temp.)(-1318.064 J) = (7.850 J/C * Final Temp.)(-196.25 J) (52.062 J/C * Final Temp.) = 1121.814 J Final Temp. = 21.547 C

OpenStudy (anonymous):

Oh, I thought you meant that the problem was you MEASURED the temperature to be lower. The calculation is incorrect..... the two MC delta T's add to zero, so one equals the negative of the other.